OpenStudy (anonymous):

Use the elimination method to solve the system of equations. Choose the correct ordered pair. 2y = x + 2 x - 3y = -5 A. (8,5) B. (6,4) C. (2,2) D. (4,3)

Parth (parthkohli):

x = 2y - 2 (2y - 2) - 3y = -5 2y - 2 - 3y = -5 2y - 3y = -3 -y = -3 y = 3 x = 4/ ANS- D(4,3)